OpenStudy (anonymous):

Elle is pushing a box of mass 7.0 kilograms with the force of 25 newtons. If the force of friction is 2.6 newtons, what is the value of acceleration of the box? A. 0.7 B. 2.6 C. 3.2 D . 3.6

OpenStudy (dream95):

F=ma 22.4=7 X a so a=3.2

OpenStudy (anonymous):

THANK YOU..I HAVE MORE DO YOU MIND HELPING?:)

OpenStudy (dream95):

you just ask

OpenStudy (anonymous):

3.2 m/s^2

OpenStudy (anonymous):

f=ma 2qn 25=7a.....(1) 2.6=ma....(2) 1qn a=25/7 2qn 2.6=ma 2.6=25/7a m=2.6x7/25 f=ma 25=2.6x7/25(a) a=625/18.2